Ellis Transport Company acquired 1.1 million shares of stock in L&K Corporation at $55 per share. They are classified by Ellis as “available for sale.” Ellis sold 100,000 shares at $59, received a 10% stock dividend, and then later in the year sold another 100,000 shares at $54.

Hint: There is no entry for the stock dividend, but a new investment per share must be calculated for use later when the shares are sold.

Required:
Prepare journal entries to record these transactions:

1) Record acquisition of shares in L & K Corporation.

2) Record sale of 100,000 shares.

3) Record sale of 100,000 shares.
Answer #1

Sl. No. Date Credit Debit 60,500,000 1 60,500,000 5900000 Description Post Ref. Investment in stocks Bank Bank Investment in

Value of Investment after 1st Sale = 60,500,000 - 5,500,000 = 55,000,000

Value of Investment after 2nd Sale = 55,000,000 - 5,500,000 = 49,500,000